OpenStudy (anonymous):

a(x-3)+8=b(x-1) Solve for x and check for any restrictions

OpenStudy (anonymous):

ax-3a+8=bx-b ax-bx=3a-8-b x(a-b) = 3a-b-8 x=(3a-b-8)/(a-b) => a must not equal b

OpenStudy (anonymous):

@ajdirig Thank you!!! :) Makes sense now

OpenStudy (anonymous):

no problem!